The sum of two consecquitive positive even numbers is 340. find the numbers
Thank you

Answer by fractalier(2101)

Let the first number be x and the next even one would be x+2.
Then by adding we get
x + (x+2) = 340
2x + 2 = 340
2x = 338
x = 169
and x+ 2 = 171
but notice these are ODD.
Since the question called for even, there is NO SOLUTION.
Tis kind of a trick question...